# High-Capacity Silicon Microdisk Transmitter

### v1

@@ III.3 Subsection 3: Silicon Oxidation
Oxidizing the silicon core is an attractive post-fabrication technique because the accompanying reduction of the effective refractive index produces a permanent blue-shift of the resonance. In our devices, however, the microrings are buried beneath a thick upper SiO2 cladding and overlaid by a metal heater. We therefore explored a two-step process—wet removal of the cladding followed by thermal oxidation of the exposed waveguide. Before the wet etching we open a window to protect the other part of the device. The window was opened using the lithography and developing. Due to the charging effect, the window was opened using the UV maskless lithography. However, even though the other part of the device is protected using the resist, the HF-based etchlant that stripped the oxide also attacked the heater electrodes, leaving the device electrically open and severely degrading the modulator (Figure 11). Owing to this collateral damage we abandoned the oxide-removal/oxidation route.

### v2

@@ III.5 Subsection 5: Silicon Oxidation
Oxidizing the silicon core is an attractive post-fabrication technique because the accompanying reduction of the effective refractive index produces a permanent blue-shift of the resonance. In our devices, however, the microrings are buried beneath a thick upper SiO2 cladding and overlaid by a metal heater wires and contact. We therefore explored a two-step process—wet removal of the cladding followed by thermal oxidation of the exposed waveguide. Before the wet etching we open a window to protect the other part of the device. The window was opened using the lithography and developing. Due to the charging effect, the window was opened using the UV maskless lithography. However, even though the other part of the device is protected using the resist, the HF-based etchlant that stripped the oxide also attacked the heater electrodes, leaving the device electrically open and severely degrading the modulator (Figure 12). Owing to this collateral damage we abandoned the oxide-removal/oxidation route.